OpenStudy (anonymous):

Find the vertex of the parabola formed by the following quadratic function: y = 3x2 – 6x + 2

OpenStudy (anonymous):

@kyusakazaki

OpenStudy (anonymous):

1, -1 i think

OpenStudy (anonymous):

@xo_ellaaavasquez x=6/6 ,y=3+2-6=5-6=-1

OpenStudy (anonymous):

yes thats correct thanks

OpenStudy (anonymous):

The vertex is at (x, y) where x = -b/2a. In this equation, b = -6, and a = 3 So the point is (1, -1)